OpenStudy (anonymous):

y(θ) = 13 + 8cos(7θ) What is the largest value of the function?

jimthompson5910 (jim_thompson5910):

Hint: cosine has the smallest value of -1 and the largest value of 1 ie, if y = cos(x), then -1 <= y <= 1

OpenStudy (anonymous):

see -1<=cos x <=1 hence -1<=cos(7θ)<=1

OpenStudy (anonymous):

-1*8<=8*cos(7θ)<=1*8 -8<=8cos(7θ)<=8 or 13-8<=8*cos(7θ)<=13+8 thus largest value is 21